OpenStudy (anonymous):

Find the product. (a2 + 2a + 1)(a + 1)

OpenStudy (anonymous):

A. a2 + 3a + 2 B. a3 + 3a2 + 3a + 1 C. a3 + 2a2 + a + 1 D. a3 + 3a2 + a + 1

OpenStudy (anonymous):

is it A?

OpenStudy (anonymous):

\[(a^2 + 2a + 1)(a + 1)\] \[=a(a^2+2a+1)+(a^2+2a+1)\] \[=a^3+2a^2+a+a^2+2a+1\] \[=a^2+3a^2+3a+1\]

OpenStudy (anonymous):

or another solution: \[a^2+2a+1=(a+1)^2\] \[(a+1)^2(a+1)=(a+1)^3=a^3+3a^2+3a+1\]